OpenStudy (reneeshumpert):

5x + y = 6 5x + 3y = -4 The y-coordinate of the solution to the system shown is _____. -5 -1 .5 1

OpenStudy (reneeshumpert):

I was thinking maybe I could substitute each number in for y , but that confused me

TheSmartOne (thesmartone):

If you solve the system of equations using elimination... \(\sf ~~~5x + y = 6\\ - \\~~~5x + 3y = -4 \\------- \) Subtract both equations.

OpenStudy (reneeshumpert):

um, I may sound stupid but how would I subtract the equations? like regular subtraction?

OpenStudy (reneeshumpert):

I got x+2y=10

TheSmartOne (thesmartone):

5x-5x = 0 Just ignore the variable and subtract the 2 numbers and at the end put back the variable.

TheSmartOne (thesmartone):

So to subtract y- 3y Think of it as 1-3 And whatever you get put 'y' right next to it.

TheSmartOne (thesmartone):

So what is \(\sf 1-3=?\)

OpenStudy (reneeshumpert):

it would be -2 .. do if 5x-5x becomes 0 I wouldnt have a x variable at all anymore ?

TheSmartOne (thesmartone):

yup the x would go away becaue 0x = 0 = nothing and sine it equals to -2 we add the y back and we get \(\sf -2y\)

OpenStudy (reneeshumpert):

so the overall new equation is -2y=10 ? or did I miss something ?

TheSmartOne (thesmartone):

What is \(\sf 6-(-4)=?\)